\begin{thebibliography}{99}

\bibitem{Agmon} C. Agmon, {\it The $L^p$ approach to the Dirichlet problem}, Ann. Scuola Norm. Sup. Pisa {\bf{13}} (1959), 405 - 448.


\bibitem{Alves1} C.O. Alves, G.M. Figueiredo, {\it Existence and multiplicity of positive solutions to a p-Laplacian equation in $\mathbb{R}^N$}, Differential Integral Equations  {\bf{19}} (2006), 143 - 162.

\bibitem{Alves2} C.O. Alves, G.M. Figueiredo, {\it On multiplicity and concentration of positive solutions
for a class of quasilinear problems with critical exponential growth in $\mathbb{R}^N$}, Journal of Differential Equations  {\bf{246}} (2009), 1288 - 1311.

\bibitem{Alves3} C.O. Alves, {\it  Multiplicity of multi-bump type nodal solutions for a class of elliptic problems in $\mathbb{R}^N$}, Topological Methods in Nonlinear Analysis  {\bf{34}} (2009), 231 - 250.

\bibitem{Alves4} C.O. Alves, J.B.M. do \'{O}, O.H. Miyagaki, {\it  On a class of singular biharmonic problems involving critical exponents}, Journal of Mathematical Analysis and Applications  {\bf{277}} (2003), 12 - 26.

\bibitem{Alves5} C.O. Alves, J.M.B. do \'{O}, O.H. Miyagaki, {\it  Nontrivial solutions for a class of semilinear biharmonic problems involving critical exponents}, Nonlinear Analysis  {\bf{46}} (2001), 121 - 133.

\bibitem{Alves6} C.O. Alves, J.M.B. do \'{O}, M.A.S. Souto, {\it  Local moutain-pass for a class of elliptic problems in $\mathbb{R}^N$ involving critical growth}, Nonlinear Analysis  {\bf{46}} (2001), 495 - 510.


\bibitem{Soares2} C.O. Alves, S.H.M. Soares, {\it  Nodal solutions for singularly perturbed equations with critical exponential growth}, Journal of Differential Equations  {\bf{234}} (2007), 464 - 484.

\bibitem{AR}  A. Ambrosetti, P.H. Rabinowitz, {\it Dual variational methods in critical point theory and applications},  J. Funct. Anal. \textbf{14} (1973), 349-381.


\bibitem{Del Pino} M. del Pino, P. Felmer, {\it Local mountain pass for semilinear elliptic problems in unbounded domains}, Calc. Var. Partial
Differential Equations {\bf{4}} (1996), 121 -
137.

\bibitem{DingTanaka} Y.H. Ding, K. Tanaka, {\it  Multiplicity of positive solutions of a nonlinear Schr\"{o}dinger equation}, Manuscripta Math.  {\bf{112}} (2003), 109 - 135.

\bibitem{Floer} A. Floer, A. Weinstein, {\it  Nonspreading wave packets for the cubic Schr\"{o}dinger equations with a bounded potential}, Journal of Functional Analysis  {\bf{69}} (1986), 397 - 408.

\bibitem{Lions} P.L. Lions, {\it  The concentration compactness principle in the calculus of variations. The locally compact case, Part 2}, Analyse Nonlin. {\bf{1}}(1984), 223 - 283.

\bibitem{Ramos} M. Ramos, {\it  Uniform estimates for the biharmonic operator in $\Bbb R^N$ and applications}, Commun. Appl. Analysis, {\bf{8}} (2009) 435 - 457.

\bibitem{Squassina} A. Salvatore, M. Squassina, {\it  Deformation from symmetry for Schr\"{o}dinger equations of higher order on unbounded domains}, Electronic Journal of Differential Equations  {\bf{65}} (2003), 1 - 15.

\bibitem{Rabinowitz} P.H. Rabinowitz, {\it  On a class of nonlinear Schr\"{o}dinger equations}, ZAMP  {\bf{43}} (1992), 270 - 291.

\bibitem{Wang} X. Wang, {\it  On concentration of positive bound states of nonlinear Schr\"{o}dinger equations}, Commum. Math. Phys.  {\bf{153}} (1993), 229 - 244.


\end{thebibliography}

\documentclass[12pt,twoside]{article}
\usepackage{amsmath,amsbsy,amsfonts,amsthm,amssymb}
\usepackage{color}
\newtheorem{lemma}{Lemma}[section]

\newtheorem{theorem}{Theorem}[section]

\newtheorem{remark}{Remark}[section]
\newtheorem*{claim}{Claim}

\let\Section=\section
\def\section{\setcounter{equation}{0}\Section}

\title{Existence and concentration of solutions for a class of biharmonic equations}

\author{Marcos T. O. Pimenta\thanks{Research Supported by CNPq - Brazil}\\
\noindent Instituto de Ci\^{e}ncias Matem\'{aticas} e de Computa\c{c}\~{a}o\\
\noindent Universidade de S\~ao Paulo \\
\noindent 13560-970, S\~ao Carlos - SP, Brazil.\\
\noindent e-mail: {\tt{mtopimenta@gmail.com}}\\
\mbox{}\\
and \\
\mbox{}\\
\noindent S\'{e}rgio H. M. Soares\footnote{Corresponding author. E-mail:~\textsf{monari@icmc.usp.br}, Phone: +55\,16\,3373\,9660, Fax: +55\,16\,3373\,9650. } \\
\noindent Departamento de Matem\'atica \\
\noindent Instituto de Ci\^{e}ncias Matem\'{a}ticas e de Computa\c{c}\~{a}o\\
\noindent Universidade de S\~ao Paulo \\
\noindent 13560-970, S\~ao Carlos - SP, Brazil. \\
\noindent e-mail: {\tt{monari@icmc.usp.br}}\\
}
\date{}

\begin{document}

\maketitle

\begin{abstract}
Some superlinear fourth order elliptic equations are considered. Ground states are
proved to exist and to concentrate at a point in the limit. The proof
relies on variational methods, where the existence and concentration
of nontrivial solutions are related to a suitable truncated equation.

\end{abstract}

{\scriptsize{\bf 2000 Mathematics Subject Classification:} 35J60,
35J35.}

{\scriptsize{\bf Keywords:} Variational methods, biharmonic equations, nontrivial solutions.}

\section{Introduction}

In the last years, many authors have been studied several questions about the following Schr\"{o}dinger elliptic equation
\begin{equation}
\left\{
\begin{array}{l}
-\epsilon^2 \Delta u + V(x)u  =  f(x,u) \ \ \mbox{in $\Omega$}\\
\ u\in H^1(\Omega)
\end{array} \right. \label{P1}
\end{equation}
with Neumann or Dirichlet boundary conditions, where $\Omega \subset \mathbb{R}^N$. Motivated by Floer and Weinstein \cite{Floer}, Rabinowitz in \cite{Rabinowitz}
uses a mountain-pass type argument to find a ground state solutions to (\ref{P1}) for $\epsilon > 0$  sufficiently small, when
$N\geq 3$, $\Omega = \mathbb{R}^N$, $f$ is a subcritical and superlinear  nonlinearity function and the potential $V$ is nonnegative and assumed to satisfy the condition
\begin{equation}\label{AR}
0 < V_0 = \inf_{x\in\mathbb{R}^N} V(x) < \liminf_{|x|\to \infty} V(x).
\end{equation}
In \cite{Wang},  Wang proves that the mountain-pass
solutions found in \cite{Rabinowitz}  concentrate around a global minimum of $V$ as $\epsilon\rightarrow 0$.
In \cite{Alves1,Alves2}, Alves and Figueiredo consider the problem (\ref{P1}) with the Laplace operator replaced by p-Laplace operator  obtaining  existence, multiplicity and concentration of positive solutions. In the celebrated paper \cite{Del Pino}, del Pino and Felmer obtained existence and concentration of solutions for the problem (\ref{P1}), where $N\geq 3$, $f$ is a subcritical and superlinear  nonlinearity function and the potential $V$ is nonnegative and it is assumed to satisfy the following condition
$$
\inf_{x\in\Lambda}V(x) < \inf_{x\in\partial\Lambda}V(x),
$$
where $\Lambda$ is a bounded domain compactly contained in $\Omega$. They developed a penalization-type
method in order to overcome the lack of compactness and used the Mountain Pass theorem to get existence and concentration of solutions. These arguments have inspired many authors in the last years, among them we could cite \cite{Alves3} and \cite{DingTanaka}, where they have obtained multiplicity and concentration of nodal and positive solutions, respectively, to an equation related to (\ref{P1}). In \cite{Soares2}, Alves and Soares obtain existence and concentration of nodal solutions of (\ref{P1}) for the case where the function $f$ has critical exponential growth.

Although there are many works dealing with problem (\ref{P1}) and with related p-Laplacian ones, just few works can be found dealing with biharmonic or even polyharmonic Schr\"{o}dinger equations. Among then we could cite \cite{Alves4} and  \cite{Alves5}, where they have obtained nontrivial solutions to semilinear biharmonic problems with critical nonlinearities and also \cite{Squassina}, where they obtained infinitely many solutions for a polyharmonic Schr\"{o}dinger equation with non-homogenous boundary data on unbounded domains.

Motivated by the results just described, a natural question is whether same
phenomenon of concentration occurs for the following class of fourth order elliptic equations
\begin{equation}
\left\{
\begin{array}{l}
\epsilon^4\Delta^2u + V(x)u = f(u) \ \ \mbox{in
$\mathbb{R}^N$}\\
u\in H^2(\mathbb{R}^N),
\end{array} \right. \label{P2}
\end{equation}
where  $\Delta^2$ is the biharmonic operator,  $\epsilon>0$ and $N\geq 5$. The functions $f:\mathbb{R}\rightarrow \mathbb{R}$ and $V:\mathbb{R}^N\rightarrow \mathbb{R}$ satisfy the following assumptions:
\begin{description}
\item [$(V_1)$] $V\in C^0(\mathbb{R}^N) \cap L^{\infty}(\mathbb{R}^N)$,
\item [$(V_2)$] There exists a bounded domain $\Omega\subset\mathbb{R}^N$ and $x_0\in\Omega$, such that $$0 < V(x_0) = V_0 = \inf_{\mathbb{R}^N}V < \inf_{\partial\Omega}V,$$
\item [$(f_1)$] $f\in C^1(\mathbb{R})$,
\item [$(f_2)$] $f(0)=f'(0)=0$,
\item [$(f_3)$] There exist constants $c_1,c_2>0$ and
$p\in\left(1,2_* - 1\right)$, such that $$|f(s)|\leq c_1 |s| +
c_2|s|^p,\quad \forall s\in\mathbb{R},
$$
 where $2_* = {2N}/{(N-4)}$,
\item [$(f_4)$] There exists $\mu > 2$ such that $0<\mu F(s)\leq sf(s)$,
for all $s \neq 0$,
\item [$(f_5)$] The function ${f(s)}/{s}$ is increasing for $s>0$ and decreasing for $s<0$.
\end{description}

Our main result is the following:
\begin{theorem}
Assume that conditions \emph{$(V_1)$}, \emph{$(V_2)$} and \emph{$(f_1) - (f_5)$} hold. Then for each sequence $\epsilon_n\rightarrow 0$, there exists a subsequence $\{\epsilon_k\}_{k\in\mathbb{N}}$ such that, for all $k\in\mathbb{N}$, there exists a nontrivial weak solution $u_k$ of (\ref{P2}) (with $\epsilon=\epsilon_k$). Moreover, if $x_k$ is the maximum point of $|u_k|$, then $x_k\in \Omega$ and $$\lim_{k\to\infty}V(x_k) = \inf_{\mathbb{R}^N} V.$$ \label{theorem1}
\end{theorem}

It is worth pointing out that although the principal arguments used can be found in \cite{Del Pino}, the approach employed here is different because the lack of a general maximum principle to the biharmonic operator. Another difficulty in dealing with biharmonic equations
   is the application of  Moser's iteration technique which would be useful in our method. In order to overcome these difficulties, we use some compactness results on Nehari manifolds, some arguments found in \cite{Alves6} and an a priori estimate for solutions of subcritical biharmonic problems found in \cite{Ramos} to prove an uniform decay of translations of $u_n$.

In the second section, we use the argument given by \cite{Del Pino} to modify the function $f$ to get the Palais-Smale condition for the functional associated with the respective modified equation. The existence and concentration of solutions to the modified problem are established.  In order to prove that this family of solutions  solves the original problem, we prove in the third section that these solutions have a kind of uniform decay at infinity.

\section{Preliminary results}{\label{variational}}

We start observing that the following problem
\begin{equation}
\left\{
\begin{array}{cl}
\Delta^2v + V(\epsilon x)v  =  f(v) & \mbox{in
$\mathbb{R}^N$}\\
v\in H^2(\mathbb{R}^N),
\end{array} \right. \label{P3}
\end{equation}
is equivalent to (\ref{P2}). In fact, the solutions $v_\epsilon$ of (\ref{P3}) and $u_\epsilon$ of (\ref{P2}) are related by
$$v_\epsilon(x) = u_\epsilon(\epsilon x).$$

Let $\mu$ be as in $(f_4)$ and let us choose $k>0$ such that $k>\frac{\mu}{\mu - 2}$. Let $a>0$ be a number such that $\max\left\{\frac{f(a)}{a},\frac{f(-a)}{-a}\right\} \leq \frac{V_0}{k}$. Set
$$
\tilde{f}(s) = \left\{
\begin{array}{lll}
\frac{-f(-a)}{a}s & \mbox{if} & s < -a\\
f(s) & \mbox{if} & |s| \leq a\\
\frac{f(a)}{a}s & \mbox{if} & s > a
\end{array}\right.
$$
and define
\begin{equation}
g(x,s) = \chi_\Omega(x)f(s) + (1-\chi_\Omega(x))\tilde{f}(s). \label{g}
\end{equation}
By $(f_1)-(f_5)$, $g$ satisfies
\begin{description}
\item [$(g_1)$] $g(x,s) = o(|s|)$ as $s\rightarrow 0$,
\item [$(g_2)$] There exist $c_1,c_2>0$ and
$p\in\left(1,2_* - 1\right)$, such that $|g(x,s)|\leq c_1 |s| +
c_2|s|^p$, for all $s\in\mathbb{R}$ and $x\in\mathbb{R}^N$,
\item [$(g_3)$] There exists $\mu>2$ such that\\
$0<\mu G(x,s)\leq g(x,s)s$,\  for all $x\in\Omega$ and $s \neq 0,$\\
$0<2 G(x,s)\leq g(x,s)s \leq \frac{1}{k}V(x)s^2$,\  for all $x\not\in\Omega$ and $s \neq 0$,
\item [$(g_4)$] ${g(x,s)}/{s}$ is nondecreasing for $s>0$ and nonincreasing for $s<0$, where $x\in\mathbb{R}^N$.
\end{description}
The problem we now consider is the following:
\begin{equation}
\left\{
\begin{array}{l}
\Delta^2v + V(\epsilon x)v = g(\epsilon x,v) \ \ \mbox{in
$\mathbb{R}^N$}\\
v\in H^2(\mathbb{R}^N).
\end{array} \right. \label{P4}
\end{equation}
Let   $E_\epsilon = \left(H^2(\mathbb{R}^N), \langle \cdot , \cdot \rangle_\epsilon \right)$  the Hilbert space endowed with the inner product
$$\langle u ,v \rangle_\epsilon = \int_{\mathbb{R}^N}\left(\Delta u \Delta v + V(\epsilon x)uv \right) dx.$$
Denote by $\|\cdot\|_\epsilon$ the norm associated with this the inner product. We consider the functional  $I_\epsilon$ defined on  $E_\epsilon$ by
$$
I_\epsilon(u) = \frac{1}{2}\int_{\mathbb{R}^N}\left(|\Delta u|^2 + V(\epsilon x)u^2\right)dx - \int_{\mathbb{R}^N}G(\epsilon x,u)dx,$$
where $G(x,s)=\int_0^s g(x,t)dt$.  The functional $I_\epsilon \in C^1(E_\epsilon, \mathbb{R}^N)$ and
\[
I'_\epsilon(u)v = \int_{\mathbb{R}^N}\left(\Delta u \Delta v + V(\epsilon x)uv \right) dx - \int_{\mathbb{R}^N}g(x,u)vdx,
\]
for all $u, v \in E_\epsilon$. Hence, critical points of $I_\epsilon$ are weak solutions of the Euler-Lagrange equation (\ref{P4}).

Our first lemma provides conditions under which $I_\epsilon$ satisfies the hypotheses of the  Mountain Pass Theorem.

\begin{lemma}\label{lemma1}
Assume that conditions $(g_1)-(g_3)$ and $(V_1)$ hold.  Then, for each $\epsilon >0$, there exist positive constants $\rho, \beta$ and $\phi\in E_\epsilon$ with $\|\phi\|_\epsilon > \rho$,  such that
\begin{enumerate}
\item $I_\epsilon(u) \geq \beta$ for all $\|u\|_\epsilon = \rho$.
\item  $I_\epsilon (\phi) <0$.
\end{enumerate}
\end{lemma}
\noindent \textbf{Proof.}
By $(g_1)$ and $(g_2)$, given $\eta>0$ there exists $A_\eta>0$ such that
$
G(x,s) \leq \eta|s|^2 + A_\eta|s|^{p+1}$
 for all $s\in \mathbb{R}$. This implies that
 $$
 \int_{\mathbb{R}^N}G(x,u)dx = o(\|u\|_\epsilon^2)$$
  as $\|u\|_\epsilon\rightarrow 0$. Then $$
  I_\epsilon(u) = \frac{1}{2}\|u\|^2 + o(\|u\|_\epsilon^2),$$ and (1) is proved.  In order to show (2), let $\varphi\in C^\infty_0(\Omega_\epsilon)$ be a nontrivial function, where $\Omega_\epsilon = \{\epsilon^{-1}x; \, x\in\Omega\}$. From $(g_3)$,
$$I_\epsilon(t\varphi) \leq \frac{t^2}{2}\|\varphi\|^2_\epsilon - Ct^\mu\int_\Omega |\varphi|^\mu dx,$$
for all $t>0$. Then $I_\epsilon(t\varphi)\rightarrow -\infty$ as $t\rightarrow +\infty$. Taking $\phi = t\varphi$ with $t>0$ sufficiently large, we obtain (2) and the proof is complete.        \quad\hfill $\Box$

\begin{lemma}\label{lemma2}
Assume that conditions $(g_1)-(g_3)$ and $(V_1)$ hold.  Then, the functional  $I_\epsilon$ satisfies the Palais-Smale condition, that is, if $\{u_n\}$ is a sequence in $E_\epsilon$ such that $\{I_\epsilon(u_n)\}$ is bounded and $I'_\epsilon(u_n) \to 0$, then $\{u_n\}$ contains a  strongly convergent subsequence in $E_\epsilon$.
\end{lemma}
\noindent \textbf{Proof.}
We start proving that  $\{u_n\}$ is a  bounded  sequence in $E_\epsilon$. From $(g_3)$ and $(V_1)$, we have
\begin{eqnarray*}
I_\epsilon(u_n) - \frac{1}{\mu}I_\epsilon'(u_n)u_n &\geq& \frac{\mu - 2}{2\mu}\|u_n\|_\epsilon^2 + \frac{1}{\mu}\int_{{\mathbb{R}^N\backslash \Omega_\epsilon}}\!\!\!\!\!\left(g(\epsilon x,u_n)u_n - \mu G(\epsilon x,u_n)\right)dx\\
& \geq & \frac{\mu - 2}{2\mu}\|u_n\|_\epsilon^2 + \frac{2-\mu}{2k\mu}\int_{\mathbb{R}^N\backslash \Omega_\epsilon}\!\!\!V(\epsilon x)u_n^2dx\\
& \geq & C \|u_n\|_\epsilon^2.
\end{eqnarray*}
Since
$$
I_\epsilon(u_n) - \frac{1}{\mu}I_\epsilon'(u_n)u_n \leq d + C\|u_n\|_\epsilon + o_n(1),
$$
for some $d\in  \mathbb{R}$, the sequence  $\{u_n\}$ is bounded on $E_\epsilon$. Hence, we can assume that
\begin{equation}
\begin{array}{ll}
u_n\rightharpoonup u & \mbox{in $E_\epsilon$},\\
u_n\rightarrow u & \mbox{in $L^q_{loc}(\mathbb{R}^N)$, for all  $1\leq q<2_*$},\\
u_n\rightarrow u & \mbox{a.e. in $\mathbb{R}^N$}
\end{array} \label{eq1}
\end{equation}
as $n\rightarrow\infty$. By Lebegue's convergence theorem, it is a simple matter to verify  that $u$ is a weak solution of (\ref{P4}).  We now take advantage of the Hilbertian structure of $E_\epsilon$ to prove that $u_n\rightarrow u$, as $n\rightarrow\infty$, by proving that $\|u_n\|_\epsilon \rightarrow \|u\|_\epsilon$ as $n\rightarrow\infty$. As we will see, this follows from the following claim.
\begin{claim}
Given $\delta>0$, there exists $R = R(\delta)>0$ such that
$$\limsup_{n\to\infty} \int_{\mathbb{R}^N\backslash B_R}\left(|\Delta u_n|^2 + V(\epsilon x)u_n^2\right)dx < \delta.
$$
\end{claim}
Effectively, for each $R>0$ let $\eta_R\in C^\infty(\mathbb{R}^N)$  be a cut-off function such that $0\leq\eta_R\leq 1$, $\eta_R = 0$ in $B_{{R}/{2}}(0)$, $\eta_R = 1$ in $B_R(0)^c$, $|\nabla\eta_R|\leq {C}/{R}$ and $|\Delta\eta_R|\leq {C}/{R^2}$. From  $(g_3)$ and H\"{o}lder inequality, for  $R>0$ such that $\Omega_\epsilon\subset B_{{R}/{2}}(0)$, we have
$$
\left(1-\frac{1}{k}\right)\int_{B_R(0)^c}\left(|\Delta u_n|^2 + V(\epsilon x)u_n^2\right)dx \leq I_\epsilon'(u_n)(\eta_Ru_n) + \frac{C}{R},
$$
and the claim follows taking the supremum limit.

Combining the claim with the Sobolev embeeding theorem and the integrability of $x\mapsto g(\epsilon x,u(x))u(x)$, we have that  given $\delta>0$ there exists $R_\delta>0$ such that
$$\limsup_{n\to\infty} \int_{B_R(0)^c}g(\epsilon x, u_n)u_n dx < \frac{\delta}{2},
$$
and
$$
\int_{B_R(0)^c}g(\epsilon x, u)u dx < \frac{\delta}{2},
$$
for all $R>R_\delta$.  By (\ref{eq1}) and $(g_2)$, we get
$$
\left|\int_{\mathbb{R}^N}(g(\epsilon x,u_n)u_n - g(\epsilon x,u)u)dx\right| < \delta,
$$
provided $n$ is  sufficiently large. Therefore,
\[
\|u_n\|_\epsilon - \|u\|_\epsilon  =  \int_{\mathbb{R}^N}(g(\epsilon x,u_n)u_n -g(\epsilon x,u)u) dx + o_n(1)
 <  \delta +o_n(1)
\]
for all $\delta>0$. The proof of Lemma {lemma2} is complete. \quad\hfill $\Box$

By the Moutain Pass Theorem \cite{AR}, for any  $\epsilon>0$  there exists $v_\epsilon \in E_\epsilon$ a weak solution of (\ref{P4}) such that $I_\epsilon(v_\epsilon) = c_\epsilon$, where
$$c_\epsilon = \inf_{\gamma\in \Gamma_\epsilon} \max_{t\in[0,1]} I_\epsilon(\gamma(t))$$
and $\Gamma_\epsilon = \{\gamma\in C^0([0,1],E_\epsilon); \,  \gamma(0)=0 \ \ \mbox{and} \ \ I_\epsilon(\gamma(1))<0\}.$

From $(g_4)$,  the minimax level $c_\epsilon$ can be characterized as (see \cite{Rabinowitz})
$$
c_\epsilon = \inf_{u\in E_\epsilon\backslash\{0\}} \max_{t\geq 0}I_\epsilon(tu) = \inf_{\mathcal{N}_\epsilon} I_\epsilon,
$$
where  $\mathcal{N}_\epsilon$ is defined by
$$\mathcal{N}_\epsilon = \left\{ u\in E_\epsilon \backslash\{0\}; \, I_\epsilon'(u)u = 0\right\}.
$$
We now consider a sequence $\{\epsilon_n\}$ with  $\epsilon_n\rightarrow 0$ as $n\rightarrow\infty$. We claim that there exists a subsequence $\{\epsilon_k\}$ such that $v_k:=v_{\epsilon_k}$ is a solution of (\ref{P3}).  The proof will be carried out by a series of lemmas.
 The first one states the existence of a ground-state solution to the limit problem.
\begin{lemma}
Suppose that $f$ satisfies $(f_1) - (f_5)$. Then, there exists a ground-state solution to the following problem
\begin{equation}
\begin{array}{cl}
\Delta^2 w + V_0w  =  f(w)& \mbox{in $\mathbb{R}^N$},
\label{P5}
\end{array}
\end{equation}
at the level
$$c_0 = \inf_{\gamma\in\Gamma_0}\sup_{0\leq t\leq 1}I_0(\gamma(t)),$$
where $I_0$ is the energy functional associated to (\ref{P5}) and $$\Gamma_0 = \{\gamma\in C^0([0,1],H^2(\mathbb{R}^N)); \,  \gamma(0)=0 \ \ \mbox{and} \ \ I_0(\gamma(1))<0\}.$$ \label{lemma3}
\end{lemma}
\noindent \textbf{Proof.} A proof can be found in \cite[Theorem 4.23]{Rabinowitz}. \quad\hfill $\Box$
\begin{lemma}
Suppose that $g$ satisfies $(g_1)-(g_4)$ and $V$ satisfies $(V_1)-(V_2)$. Then
$$
\limsup_{n \to \infty} c_{\epsilon_n} \leq c_0.
$$
\label{lemma4}
\end{lemma}
\noindent \textbf{Proof.}
We assume without loss of generality that $x_0 = 0$, for $x_0$ given by condition $(V_2)$.  Let $w$ be a solution of (\ref{P5}) such that $I_0(w)=c_0$.
Let $\psi\in C^\infty(\mathbb{R}^N)$ be a cut-off function such that $\psi \equiv 1$ in $B_\rho(0)$ and $\psi \equiv 0$ in $B_{2\rho}^c(0)$, where $B_{2\rho}(0)\subset \Omega$. Set $w_n(x) = \psi(\epsilon_n x)w(x)$ and note that $supp (w_n) \subset \Omega_{\epsilon_n}$, $w_n\rightarrow w$ in $H^2(\mathbb{R}^N)$ and in $L^r(\mathbb{R}^N)$ where $2\leq r \leq 2_*$. Let $\varphi_{\epsilon_n}(w_n) > 0$ be such that $\varphi_{\epsilon_n}(w_n)w_n\in \mathcal{N}_{\epsilon_n}$.
An easy computation shows that $\varphi_{\epsilon_n}(w_n) \rightarrow 1$ as $n\rightarrow \infty$. Then
\begin{eqnarray*}
c_{\epsilon_n} & \leq & I_{\epsilon_n}(\varphi_{\epsilon_n}(w_n)w_n)\\
& = & I_0(\varphi_{\epsilon_n}(w_n)w_n) + \frac{1}{2}\int_{\mathbb{R}^N}\left(V(\epsilon_n x) - V(0)\right)(\varphi_{\epsilon_n}(w_n)w_n)^2 dx,
\end{eqnarray*}
and the result follows by the Lebesgue Dominated Convergence Theorem.
\quad\hfill $\Box$

It is easy to see from the last lemma that $\{v_n\}$ is a bounded sequence in $H^2(\mathbb{R}^N)$.

\begin{lemma}
There exists $\{y_n\}\subset \mathbb{R}^N$ and $R,\beta >0$ such that
$$\liminf_{n \to \infty} \int_{B_R(y_n)} v_n^2 dx \geq \beta > 0.$$ \label{lemma5}
\end{lemma}
\noindent \textbf{Proof.}
Suppose the assertion of the lemma is false. Then by Lemma I.1 of \cite{Lions} (with $q=2$ and $p=\frac{2N}{N-2}$ ), $v_n\rightarrow 0$ in $L^r(\mathbb{R}^N)$ where $2\leq r \leq 2_*$. Hence by the Lebesgue Dominated Convergence Theorem, we get
$$\int_{\mathbb{R}^N} g(\epsilon_n x, v_n) v_n dx = o_n(1) \,\ \mbox{and} \,\ \int_{\mathbb{R}^N} G(\epsilon_n x, v_n)dx = o_n(1).$$
Then $c_{\epsilon_n}\rightarrow 0$ as $n\rightarrow \infty$. On the other hand, since the minimax value is an increasing function of the potential we have $c_{\epsilon_n} \geq d$, $\forall n\in \mathbb{N}$, where $d>0$ is the minimax value associated to the problem
$$
\Delta^2 v + V_0 v = g(\epsilon_n x, v) \,\ \mbox{in $\mathbb{R}^N$}.
$$ This contradiction proves the lemma.
\quad\hfill $\Box$

For $R>0$ given by Lemma \ref{lemma5}, we have:
\begin{lemma}
The sequence $\{\epsilon_n y_{\epsilon_n}\}$ is bounded and
$dist(\epsilon_n y_{\epsilon_n},\Omega) \leq \epsilon_n R$.   \label{lemma6}
\end{lemma}
\noindent \textbf{Proof.}
Let $K_\delta$  denote a $\delta$-neighborhood of $\Omega$, where $\delta>0$. Let $\phi\in C^\infty(\mathbb{R}^N)$ be a cut-off function such that $\phi = 0$ in $\Omega$, $\phi = 1$ in $\mathbb{R}^N\backslash K_\delta$, $0 \leq \phi \leq 1$, $|\nabla\phi| \leq {C}/{\delta}$ and $|\Delta\phi| \leq {C}/{\delta^2}$. Setting $\phi_\epsilon(x) = \phi(\epsilon x)$ and using $v_{\epsilon_n}\phi_{\epsilon_n}$ as test function in (\ref{P4}) we have
$$\int_{\mathbb{R}^N}\left(\Delta v_{\epsilon_n}\Delta(v_{\epsilon_n}\phi_{\epsilon_n}) + V(\epsilon_n x)v_{\epsilon_n}^2\phi_{\epsilon_n}\right)dx = \int_{\mathbb{R}^N}g(\epsilon_n x, v_{\epsilon_n})v_{\epsilon_n}\phi_{\epsilon_n} dx,$$
which gives
\begin{eqnarray*}
V_0\left(1-\frac{1}{k}\right)\int_{\mathbb{R}^N}v_{\epsilon_n}^2 \phi_{\epsilon_n} dx & \leq &
\frac{C\epsilon_n}{\delta}\|v_{\epsilon_n}\|^2_{H^2(\mathbb{R}^N)} + \frac{C\epsilon_n^2}{\delta^2}\|v_{\epsilon_n}\|^2_{H^2(\mathbb{R}^N)}\\
& \leq & \frac{C\epsilon_n}{\delta}\|v_{\epsilon_n}\|^2_{H^2(\mathbb{R}^N)}.
\end{eqnarray*}
If there is $\{\epsilon_k\}$ subsequence such that $B_R(y_{\epsilon_k}) \cap {K_\delta}/{\epsilon_k} = \emptyset$, then
\begin{eqnarray*}
V_0\left(1-\frac{1}{k}\right)\int_{B_R(y_{\epsilon_k})}v_{\epsilon_k}^2 dx & \leq & V_0\left(1-\frac{1}{k}\right)\int_{\mathbb{R}^N}v_{\epsilon_k}^2 \phi_{\epsilon_k} dx\\
& \leq & \frac{C\epsilon_k}{\delta}\|v_{\epsilon_k}\|^2_{H^2(\mathbb{R}^N)} \rightarrow 0
\end{eqnarray*}
as $k\rightarrow \infty$, which contradicts Lemma \ref{lemma5}.  Hence, for each $n\in\mathbb{N}$ there exists $x_n$ such that $\epsilon_n x_n\in K_\delta$ and $|y_{\epsilon_n} - x_n| < R$. Hence,  $dist(\epsilon_n y_{\epsilon_n}, \Omega) < \epsilon_n R + \delta$, for all $\delta>0$,  which completes the proof.
\quad\hfill $\Box$

\begin{remark}
It is worth pointing out that by Lemma \ref{lemma6}, we can assume that $\epsilon_n y_{\epsilon_n} \in \overline{\Omega}$ for all $n$ sufficiently large. In fact, on the contrary,  we consider ${\epsilon_n^{-1}}{z_n}$  instead of  $y_{\epsilon_n}$, where $z_n\in \overline{\Omega}$ is such that $|\epsilon_n y_n - z_n| < \epsilon_n R$. This fact will be used to guarantee that $\epsilon_n y_{\epsilon_n} \rightarrow x'_0 \in \overline{\Omega}$.
\end{remark}

The following result plays a central role in the proof of Theorem \ref{theorem1}.

\begin{lemma}\label{lemma7}
The following assertions hold:
\begin{description}
\item [\emph{(i)}] $ \displaystyle \lim_{n \to \infty} c_{\epsilon_n} = c_0$,
\item [\emph{(ii)}] $ \displaystyle  \lim_{n \to \infty} V(\epsilon_n y_{\epsilon_n}) = V_0$.
\end{description}
\end{lemma}
\noindent \textbf{Proof.}
By Lemma \ref{lemma6}, we can assume that $\epsilon_n y_{\epsilon_n} \rightarrow x_0' \in \overline{\Omega}$ along a subsequence. Let us consider $w_n(x) = v_{\epsilon_n}(x+y_{\epsilon_n})$. Note that by Lemma \ref{lemma5}
$$\liminf_{n \to \infty} \int_{B_R(0)}w_n^2dx \geq \beta > 0.$$
Using that $\{v_{\epsilon_n}\}$  is bounded, it follows that there exists $w\in H^2(\mathbb{R}^N)\backslash\{0\}$, such that
$$
\begin{array}{ll}
w_n \rightharpoonup w & \mbox{in $H^2(\mathbb{R}^N)$}\\
w_n \rightarrow w & \mbox{in $L^r_{loc}(\mathbb{R}^N)$ where $2\leq r < 2_*$}\\
w_n \rightarrow w & \mbox{a.e. in $\mathbb{R}^N$}.\\
\end{array}
$$
Since  $w_n$ satisfies
\begin{equation}
\Delta^2 w_n + V(\epsilon_n x + \epsilon_n y_n)w_n = g(\epsilon_n x + \epsilon_n y_n, w_n) \,\ \mbox{in $\mathbb{R}^N$}, \label{P6}
\end{equation}
we have that $w$ satisfies
\begin{equation}
\Delta^2 w + V(x_0')w = \chi(x)f(w) + \left(1-\chi(x)\right)\tilde{f}(x) = \tilde{g}(x,w) \,\ \mbox{in $\mathbb{R}^N$}, \label{P7}
\end{equation}
where $\chi(x) = \lim_{n\to\infty}\chi_{\Omega}(\epsilon_n x+ \epsilon_n y_n)$, almost everywhere  in $\mathbb{R}^N$.

Denote by $\tilde{I}$ the energy functional associated with the problem (\ref{P7}) and by  $\tilde{c}$ its  minimax level.  We now consider the problem
\begin{equation}
\Delta^2 w + V(x_0')w = f(w) \,\ \mbox{in $\mathbb{R}^N$} \label{P8}
\end{equation}
and let denote $\bar{c}$ the minimax level of the functional $\bar{I}$ associated with  the problem (\ref{P8}).  In the following we  show that
$\bar{c} = \tilde{c}.$  Since
$$
\tilde{G}(x,s) = \int_0^s\tilde{g}(x,t)dt \leq F(s),
$$
we have $\bar{I}(u) \leq \tilde{I}(u)$ for all $u\in H^2(\mathbb{R}^N)$, which implies that $\bar{c} \leq \tilde{c}$. In order to verify that   $\tilde{c}  \leq  \bar{c}$, it is sufficient to prove that
\begin{equation}\label{upper}
\tilde{I}(w) \leq \liminf_{n\to\infty} c_{\epsilon_n}.
\end{equation}
In fact, if (\ref{upper}) holds, then
\begin{equation}
\tilde{c} \leq \tilde{I}(w) \leq \liminf_{n\to\infty} c_{\epsilon_n} \leq \limsup_{n\to\infty} c_{\epsilon_n} \leq c_0 \leq \bar{c}. \label{*}
\end{equation}
Therefore, $\tilde{c} = \bar{c}$,  which implies that
$\tilde{c} = \bar{c} = c_0.$  By (\ref{*}),  we have
$$\lim_{n\to\infty}c_{\epsilon_n} = c_0,$$ which proves \emph{(i)}.

The proof of (\ref{upper}) is based on  some ideas of \cite[Lemma 2.2]{Del Pino}. By elliptic estimates we see that $w_n$ converges to $w$ in $C^2_{loc}(\mathbb{R}^N)$. Hence, for all $R>0$, we have
\begin{eqnarray*}
\lim_{n\to\infty}\left[\int_{B_R(0)}\left(\frac{1}{2} \left(|\Delta w_n|^2 + V(\epsilon_nx + \epsilon_n y_n)w_n^2\right)- G(\epsilon_nx + \epsilon_n y_n, w_n)\right)dx\right]\\
=\int_{B_R(0)}\left(\frac{1}{2} \left(|\Delta w|^2 + V(x_0')w^2\right) - \tilde{G}(x,w)\right)dx.
\end{eqnarray*}
Since the last integral converges to $\tilde{I}(w)$ as $R\rightarrow +\infty$, for all $\delta>0$ there exists $R>0$ sufficiently large such that
\begin{eqnarray*}
&&\lim_{n\to\infty}\left[\int_{B_R(0)}\left(\frac{1}{2}\left(|\Delta w_n|^2 + V(\epsilon_nx + \epsilon_n y_n)w_n^2\right)-G(\epsilon_nx + \epsilon_n y_n, w_n)\right)dx\right] \nonumber\\
&&  \geq  \tilde{I}(w) - \delta.
\end{eqnarray*}
Hence, in order to proof (\ref{upper}), it suffices to show that for $R>0$ suffciently large, we have
\begin{eqnarray} \label{**}
&&{\liminf_{n\to\infty}\left[\int_{B_R^C(0)}\left(\frac{1}{2}\left(|\Delta w_n|^2 + V(\epsilon_nx + \epsilon_n y_n)w_n^2\right) - G(\epsilon_nx + \epsilon_n y_n, w_n)\right)dx\right] }\nonumber\\
&& \geq  -\delta.
\end{eqnarray}
Consider a smooth cut-off function $\eta_R$ such that $\eta_R = 1$ in $B_R^c(0)$, $\eta_R = 0$ in $B_{R-1}(0)$ and $|\nabla\eta_R|,|\Delta\eta_R| \leq C$, where $C$ is independent of $R$. Using  $\varphi = \eta_R w_n$ as a test function in (\ref{P6}),  yields
\begin{eqnarray*}
0 = \int_{\mathbb{R}^N}\left(\Delta w_n \Delta(\eta_R w_n) + V(\epsilon_nx + \epsilon_n y_n)\eta_Rw_n^2 - g(\epsilon_nx + \epsilon_n y_n, w_n)\eta_R w_n \right)dx\\
=  A_{1,n} + A_{2,n} + A_{3,n},
\end{eqnarray*}
where
\begin{eqnarray*}
A_{1,n} &=& \int_{B_R^c(0)}\left(|\Delta w_n|^2 + V(\epsilon_nx + \epsilon_n y_n)w_n^2 - 2G(\epsilon_nx + \epsilon_n y_n, w_n)\right)dx,\\
A_{2,n}& =& \int_{B_R^c(0)}\left(2G(\epsilon_nx + \epsilon_n y_n, w_n) - g(\epsilon_nx + \epsilon_n y_n, w_n)w_n \right)dx
\end{eqnarray*}
and
\begin{eqnarray*}
\lefteqn{A_{3,n} =}\\
&&\int_{_{A_{R,R-1}}}\!\!\!\!\!\!\!\!\!\!\!\!\left[\Delta w_n \Delta(\eta_R w_n) + V(\epsilon_nx + \epsilon_n y_n)(\eta_Rw_n)^2 - g(\epsilon_nx + \epsilon_n y_n, w_n)\eta_R w_n \right]dx
\end{eqnarray*}
where $A_{R,R-1}$ is the annulus $B_R(0)\backslash B_{R-1}(0)$.   By $(g_3)$, $A_{2,n} \leq 0$. On the other hand, we can choose $R>0$ sufficiently large such that
$$\lim_{n\to\infty}|A_{3,n}| \leq \delta.$$
Hence, as $A_{1,n} = -A_{2,n} -A_{3,n}$,
$$
\liminf_{n\to\infty}\frac{1}{2} A_{1,n} \geq -\frac{\delta}{2}
$$
which shows that (\ref{**}) holds. Consequently (\ref{upper}) holds.

Finally, suppose by contradiction that \emph{(ii)} is false.  Thus,  $V(x'_0) > V_0$ and so $\tilde{c}>c_0$, which is impossible. This concludes the proof of Lemma \ref{lemma7}. \quad\hfill $\Box$

\medskip
By the proof of Lemma \ref{lemma7}, we see that $\epsilon_n y_n \rightarrow x_0'\in\overline{\Omega}$. Nevertheless, by $(V_2)$  it follows that $x'_0\in \Omega$, which implies that $\chi \equiv 1$ and that $w$ satisfies (\ref{P5}). This fact will be used in the final argument of the proof of Theorem \ref{theorem1}, in the next section.

\section{Uniform decay}

Although we have obtained existence and concentration of solutions to the modified problem, nothing can be said about the original one. In order to prove that the function $v_n$ is in fact a solution of the original problem, we will prove a kind of uniform decay at infinity of the translations $w_n$.  We first establish a compactness result.

\begin{lemma}\label{compact}
Let $\{z_n\}\subset H^2(\mathbb{R}^N)$ be such that $I_0(z_n)\rightarrow c_0$ and $z_n\in\mathcal{N}_0$, for all $n\in\mathbb{N}$. If $z_n\rightharpoonup z\neq 0$, then $z_n\rightarrow z$ in $H^2(\mathbb{R}^N)$ along a subsequence. \label{lemma9}
\end{lemma}
\noindent \textbf{Proof.}
By the Ekeland Variational Principle,  there exists a sequence $\{\hat{z}_n\}\subset \mathcal{N}_0$  such that
\[
I_0(\hat{z}_n)\rightarrow c_0, \quad  I_0'(\hat{z}_n)\rightarrow 0\quad \mbox{and}\quad  \|\hat{z}_n - z_n\|_{H^2(\mathbb{R}^N)} < \frac{1}{n}.
\]
It is easy to verify that $\{\hat{z}_n\}$ is a bounded sequence in $H^2(\mathbb{R}^N)$. Thus along a subsequence $\hat{z}_n \rightharpoonup z, \ \ \mbox{in $H^2(\mathbb{R}^N)$.}$   In the following, we  prove that $\|\hat{z}_n\|_{H^2(\mathbb{R}^N)} \rightarrow \|z\|_{H^2(\mathbb{R}^N)}$ as $n\rightarrow \infty$.
Since  $z$ is a weak solution of (\ref{P5}),  we have
\begin{eqnarray*}
\int_{\mathbb{R}^N}\left(|\Delta z|^2 + V_0z^2\right)dx & \leq & \liminf_{n \to \infty} \int_{\mathbb{R}^N}\left(|\Delta \hat{z}_n|^2 + V_0\hat{z}_n^2\right)dx\\
& \leq & \limsup_{n \to \infty} \int_{\mathbb{R}^N}\left(|\Delta \hat{z}_n|^2 + V_0\hat{z}_n^2\right)dx\\
& \leq & \int_{\mathbb{R}^N}\left(|\Delta z|^2 + V_0z^2\right)dx.
\end{eqnarray*}
To verify the last inequality, assume that $$\int_{\mathbb{R}^N}\left(|\Delta z|^2 + V_0z^2\right)dx < \limsup_{n \to \infty} \int_{\mathbb{R}^N}\left(|\Delta \hat{z}_n|^2 + V_0\hat{z}_n^2\right)dx.$$ Then by Fatou's Lemma
\begin{eqnarray*}
c_0 & = & \lim_{n\to\infty} I_0(\hat{z}_n)
= \lim_{n\to\infty} \left( I_0(\hat{z}_n) - \frac{1}{\mu}I_0'(\hat{z}_n)\hat{z_n}\right)\\
& = & \limsup_{n\to\infty}\left(\frac{1}{2} - \frac{1}{\mu}\right)\int_{\mathbb{R}^N}\left(|\Delta \hat{z}_n|^2 + V_0\hat{z}_n^2\right)dx\\
& & +\  \liminf_{n\to\infty}\int_{\mathbb{R}^N}\left(\frac{1}{\mu}f(\hat{z}_n)\hat{z}_n - F(\hat{z}_n)\right)dx\\
& > & \left(\frac{1}{2} - \frac{1}{\mu}\right)\int_{\mathbb{R}^N}\left(|\Delta z|^2 + V_0z^2\right)dx + \int_{\mathbb{R}^N}\left(\frac{1}{\mu}f(z)z - F(z)\right)dx\\
& = & I_0(z) \geq c_0.
\end{eqnarray*}
This contradiction proves the result.\quad\hfill $\Box$

\begin{lemma} The sequence $\{w_n\}$ contains a  strongly convergent subsequence in $H^2(\mathbb{R}^N)$.  \label{lemma8}
\end{lemma}
\noindent \textbf{Proof.}
By Lema \ref{lemma7}, we have
$$
\lim_{n\to\infty} I_{\epsilon_n}(v_n) = \lim_{n\to\infty} c_{\epsilon_n} = c_0.
$$
Denote by $\mathcal{N}_0$ the Nehari manifold associated to (\ref{P5}). Given $v\in H^2(\mathbb{R}^N)\backslash\{0\}$, from $(g_4)$,  there exists  $\varphi_0(v)>0$ such that $\varphi_0(v)v \in \mathcal{N}_0$. Set $\tilde{w}_n = \varphi_0(w_n)w_n$. Hence,
\begin{eqnarray*}
c_0 & \leq & \frac{1}{2}\int_{\mathbb{R}^N}\left(|\Delta\tilde{w}_n|^2 + V_0\tilde{w}_n^2\right) dx - \int_{\mathbb{R}^N}F(\tilde{w}_n)dx\\
& \leq & \frac{1}{2}\int_{\mathbb{R}^N}\left(|\Delta\tilde{w}_n|^2 + V(\epsilon_n x + \epsilon_n y_n)\tilde{w}_n^2\right) dx - \int_{\mathbb{R}^N}G(\epsilon_n x + \epsilon_n y_n, \tilde{w}_n)dx\\
& = & I_{\epsilon_n}(\varphi_0(w_n)v_n)
\leq I_{\epsilon_n}(v_n)
 = c_{\epsilon_n} = c_0 + o_n(1),
\end{eqnarray*}
which implies that $I_0(\tilde{w}_n)\rightarrow c_0$ as $n\rightarrow \infty$.

We now prove that $\varphi_0(w_n) \rightarrow \varphi_0 > 0$ along a subsequence. We  first observe  that there exists $M>0$ such that $|\varphi_0(w_n)| \leq M$, $\forall n\in\mathbb{N}$. In fact, since $w_n \nrightarrow 0$ there exists $\delta>0$ such that $\|w_n\|_{H^2(\mathbb{R}^N)} > \delta$ along a subsequence. On the other hand,  it is easy to see that $\{\tilde{w}_n\}$ is a bounded sequence in $H^2(\mathbb{R}^N)$. Then
$$|\varphi_0(w_n)|\delta < \|\varphi_0(w_n)w_n\|_{H^2(\mathbb{R}^N)} \leq K$$
which implies that
$$|\varphi_0(w_n)| \leq \frac{K}{\delta} = M, \ \ \forall n\in\mathbb{N}.$$
Hence,  $\varphi_0(w_n)\rightarrow \varphi_0 \geq 0$.  We now observe that $\varphi_0 > 0$, otherwise
$$\|\tilde{w}_n\|_{H^2(\mathbb{R}^N)} = |\varphi_0(w_n)|\|w_n\|_{H^2(\mathbb{R}^N)} \rightarrow 0$$ as $n\rightarrow \infty$, which is impossible.  Therefore $\tilde{w}_n = \varphi_0(w_n)w_n \rightharpoonup\varphi_0 w \neq 0$ in $H^2(\mathbb{R}^N)$.  Therefore, we conclude from Lemma \ref{compact} that the lemma is proved. \quad\hfill $\Box$

\medskip
Combing  Lemma \ref{lemma8} with the Sobolev imbeddings , it follows that $w_n \rightarrow w$ in $L^{2_*}(\mathbb{R}^N)$. Therefore,  we obtain
\begin{equation}
\int_{B_R^c(0)}|w_n|^{2_*}dx \rightarrow 0 \quad \mbox{as $R\rightarrow \infty$ uniformly in $n$}.
\label{eq3}
\end{equation}

\begin{lemma}
$w_n(x) \rightarrow 0$ as $|x|\rightarrow \infty$, uniformly in $n$. \label{lemma10}
\end{lemma}
\begin{proof}
By the uniform $L^\infty$ estimates to solutions of subcritical biharmonic equations given in \cite{Ramos}, we have
$$\|w_n\|_{L^\infty(\mathbb{R}^N)} \leq C, \quad \forall\, n\in\mathbb{N},$$ where $C$ is independent of $n$.
Given any $x\in\mathbb{R}^N$, the function $w_n \in L^q(B_1(x))$ for all $q\geq 1$. By \cite[Theorem 7.1]{Agmon} it follows that
\begin{eqnarray*}
\|w_n\|_{W^{4,q}(B_1(x))} & \leq & C\left( \|f(w_n)\|_{L^q(B_2(x))} + \|w_n\|_{L^q(B_2(x))}\right) \\
& \leq & C \|w_k\|_{L^q(B_2(x))}\\
& \leq & C \|w_k\|_{L^\infty(\mathbb{R}^N)}^{\frac{q-2_*}{q}} \|w_k\|_{L^{2_*}(B_2(x))}^{2_*}\\
& = & C \|w_k\|_{L^{2_*(B_2(x))}}^{2_*},
\end{eqnarray*}
with $C>0$ being a constant independent of $x$ and $n$.   If $q>N$, we have the continuous imbedding $W^{4,q}(B_1(x))\hookrightarrow C^{3,\alpha}(\overline{B_1(x)})$ for $\alpha \in \left(0,1-\frac{N}{q}\right)$. Then
\begin{eqnarray*}
\|w_k\|_{C^{3,\alpha}(\overline{B_1(x)})} & \leq & \|w_k\|_{W^{4,q}(B_1(x))}
 \leq C \|w_k\|_{L^{2_*}(B_2(x))}^{2_*}.
\end{eqnarray*}
By (\ref{eq3}), it follows that
$|w_n(x)| \rightarrow 0$   as $|x|\rightarrow \infty$ uniformly in $n$.
\end{proof}

Finally, we are ready to  prove that $v_n$ is in fact a solution of (\ref{P3}). Let $n_0\in\mathbb{N}$ and $\rho>0$ be such that
$$|w_n(x)| < a, \quad \forall\,  x\in B_{\rho}(0)^c, \ \forall\,  n\geq n_0.$$
Since $x_0'\in\Omega$ and $\epsilon_ny_n\rightarrow x_0'$, it is possible to choose $n_1\in\mathbb{N}$ such that $B_\rho(0) \subset \left(\Omega \epsilon_n^{-1}\ - y_n\right)$, for all $n\geq n_1$. Taking $n\geq \max\{n_0,n_1\}$, we have
$$g(\epsilon_n x+\epsilon_n y_n,w_n(x)) = f(w_n(x)),\quad \forall x\in\mathbb{R}^N.$$
Hence,  for $n\geq \max\{n_0,n_1\}$ it follows that $w_n$ satisfies
$$
\Delta^2w_n + V(\epsilon_n x + \epsilon_n y_n)w_n =
f(w_n) \,\ \mbox{in $\mathbb{R}^N$},
$$
which implies that $v_n$ satisfies (\ref{P3}).

In order to prove the concentration  behavior  of solutions, we claim that there exists $\rho>0$ such that $\|u_n\|_{L^\infty(\mathbb{R}^N)} = \|w_n\|_{L^\infty(\mathbb{R}^N)} > \rho$, for all $n\in\mathbb{N}$ along a subsequence. In fact, if $\|w_n\|_{L^\infty(\mathbb{R}^N)} \rightarrow 0$, then
\begin{eqnarray*}
\|w_n\|^2_{H^2(\mathbb{R}^N)} & \leq & C\int_{\mathbb{R}^N}\left(|\Delta w_n|^2 + V(\epsilon_n x + \epsilon_n y_n)w_n^2\right)dx\\
& = & C\int_{\mathbb{R}^N}f(w_k)w_kdx\\
& \leq & \|w_k\|_{L^\infty(\mathbb{R}^N)}\int_{\mathbb{R}^N}f(w_k)dx \rightarrow 0
\end{eqnarray*}
as $n\rightarrow\infty$, which contradicts the fact that $w_n\rightarrow w$ and $w \neq 0$.

Let $x_n$ be the maximum point of $|u_n|$ in $\mathbb{R}^N$, then
$$
p_n := \frac{x_n - \epsilon_ny_n}{\epsilon_n}
$$
is the maximum point of $|w_n|$. By Lemma \ref{lemma10}, there exists $R_0 > 0$ such that $p_n \in B_{R_0}(0)$ for all $n$ sufficiently large. Then, along a subsequence $p_n \rightarrow p_0$ as $n\rightarrow \infty$. Hence
$$
x_n = \epsilon_n p_n + \epsilon_n y_n \rightarrow x_0' \in \Omega, \ \ \mbox{as $n\rightarrow \infty$},
$$
where $V(x'_0) = V_0$, which proves Theorem \ref{theorem1}.

\vspace{0.5cm}

\noindent \textbf{Acknowledgment.}\  The authors are grateful to  Profs. Claudianor O. Alves and Marco A. S. Souto for valuable discussions.

\end{document}